OpenStudy (anonymous):

8 + 3[(10 ÷ 2 ) + 6 ] P10÷2=5+6=11 E N/A M N/A D N/A A 8+3=11+11=22 S N/A is this correct?

OpenStudy (anonymous):

come again?

OpenStudy (anonymous):

8 + 3[(10 ÷ 2 ) + 6 ] P10÷2=5+6=11 E N/A M N/A D N/A A 8+3=11+11=22 S N/A is this correct?